1.a) Sketch the graphs on the same axis.
(i) y = x^2 - 4
(ii) y = -x + 3

(b) What are the inequalities of the region bounded (enclosed area) by the two graphs.
1
Expert's answer
2019-08-16T03:46:57-0400

a) "y=x^2-4" - red line;

"y=-x+3" - blue line.



b) "x^2-4=-x+3"

"x^2+x-7=0"

"x=\frac{-1-\sqrt{29}}{2}\;\;or\;\;x=\frac{-1+\sqrt{29}}{2}"

The region bounded by the two graphs: "\frac{-1-\sqrt{29}}{2}<x<\frac{-1+\sqrt{29}}{2}"

"x^2-4<y<-x+3"